The length of a rectangular playing field is 5 ft less than twice its width.
If the perimeter of the playing field is 230 ft, find the length and width of the field.

Answer by uknouluvme89(30)

Let x = the width
Let 2x-5 = length
P= 2l + 2w
230 = 2(2x-5) + 2x
230 = 4x-10+2x
230 = 6x-10
add 10 to both sides of the equation
240 = 6x
divide both sides by 6
40 = x
Width = x = 40ft
Length = 2x-5 = 2(40)-5 = 75 ft
Hope that helped!